OpenStudy (anonymous):

Match the scale factor of the side lengths of a regular polygon with the scale factor that results for the area, volume or perimeter.

OpenStudy (anonymous):

Picture

OpenStudy (anonymous):

Thats the possible answers.

OpenStudy (anonymous):

I need help its my final exam and i dont know the answer

OpenStudy (anonymous):

??

OpenStudy (cwrw238):

this question confuses me a little but if i understand it corret;y if side = 6 then volume scale factor = 6*3 = 216 for side 12 area factor = 12^2 = 144

OpenStudy (anonymous):

ya i got that set up but i guess i just did the multiplication wrong.

OpenStudy (anonymous):

but this one multiple choice so i dont understand which pieces are connected to the problem??

OpenStudy (anonymous):

Like which one is abc or d?

OpenStudy (anonymous):

What about this problem can you give me the answer to it??

OpenStudy (cwrw238):

6 = A 12 = C

OpenStudy (cwrw238):

30 = D 36 = B

OpenStudy (anonymous):

Okay thanks

OpenStudy (cwrw238):

yw gotta go now

OpenStudy (anonymous):

okay thanks for the help